# Phase transitions in an expanding medium – hot remnants

### v1

@@ Abstract
We employ a relatively simple model combining hydrodynamics and a dynamical order parameter or “mixing fraction” introduced in [1]. The model was fitted to describe the main features of the holographic Witten model [2], that were worked out in [3], with regard to the physics associated to the phase transitions. However, the structure of the model is very general and we expect that it can be applied in many other settings. The confinement-deconfinement phase transition has been studied intensively in the framework of holography (see for example [4, 5, 6, 7, 8, 9]).
@@
The main result of this letter in part challenges this simple picture: we observe the very generic appearance of hot remnants of the high energy phase, whose temperature remains in the vicinity of $T_{c}$ despite the expansion of the medium. For this reason, as the plasma is not supercooled, these remnants do not dissipate through bubble nucleation of the low energy phase, but rather shrink. Then the system starts to “dissolve” with the remaining decreasing fraction of the deconfined phase being counterintuitively even heated up. These findings seem very generic and robust and appear in such diverse scenarios as boost-invariant expansion in flat Minkowski space or cosmological expansion in a FRW universe.
The physics behind this behaviour is an interplay between the hydrodynamics of the high energy phase, latent heat and the dynamics of the order parameter. After introducing the details of our model, we will describe the various stages in the evolution of the remnants, namely the initial “heating up”, a shrinking stage and subsequent dissolution.
@@
Simplified model of the holographic system. As the bulk numerical relativity description of time dependent dynamics in the holographic Witten model is extremely challenging [10], in [1], we introduced a boundary model which describes very well the domain wall in the dual field theory, obtained through numerical relativity [3]. The model combines together the hydrodynamic description of the deconfined phase with a simple description of the confined phase, by adding an additional dynamical degree of freedom $\gamma$ describing the mixing of the two phases. Specifically, the total energy-momentum tensor of the theory is a mixture of the hydrodynamic tensor of the deconfined phase and the trivial $T^{\mu\nu}=\eta^{\mu\nu}$ of the confined phase, together with a contribution describing the domain wall ($T^{\mu\nu}_{\Sigma}$):
@@
Boost invariant expansion. Boost invariant expansion, first introduced by Bjorken [15], can be seen as a way of modelling extremely energetic heavy-ion collisions, where the system undergoes expansion along the longitudinal (collision) axis. Consequently, this leads to decreasing energy density and hence also decreasing temperature. The expansion then automatically drives the cooling system through any phase transitions or crossover to the low energy confined phase. Technically, one writes the standard Minkowski metric in adapted coordinates
@@
Neglecting the explicit temporal derivatives, the spatial part can be translated into a virial equation, but modulated by the scale factor $a(t)$. This indicates that the width of the dissolving blob shrinks with $a(t)$ in contrast to staying constant in the boost-invariant case. The virial theorem also leads to the conclusion of a rise of the temperature of the dissolving fluid. This is in very good agreement with the numerical simulations as shown in Fig. 9.

### v2

@@ Abstract
We employ a relatively simple model combining hydrodynamics and a dynamical order parameter or “mixing fraction” introduced in [1]. The model was fitted to describe the main features of the holographic Witten model [2] (without flavours), that were worked out in [3], with regard to the physics associated to the phase transitions. However, the structure of the model is very general and we expect that it can be applied in many other settings. Note that the confinement-deconfinement phase transition [4, 5, 6, 7, 8, 9] has been studied intensively using holography.
@@
The main result of this letter in part challenges this simple picture: we observe the very generic appearance of hot remnants of the high energy phase, whose temperature remains in the vicinity of $T_{c}$ despite the expansion of the medium. For this reason, as the plasma is not supercooled, these remnants do not dissipate through bubble nucleation of the low energy phase, but rather shrink. Then the system starts to “dissolve” with the remaining decreasing fraction of the deconfined phase being counterintuitively even heated up. These findings seem very generic and robust and appear in such diverse scenarios as boost-invariant expansion in flat Minkowski space or cosmological expansion in a FRW universe. The physics behind this behaviour is an interplay between the hydrodynamics of the high energy phase, latent heat and the dynamics of the order parameter. After introducing the details of our model, we will describe the various stages in the evolution of the remnants, namely the initial “heating up”, a shrinking stage and subsequent dissolution.
@@
Simplified model of the holographic system. As the bulk numerical relativity description of time dependent dynamics in the holographic Witten model is extremely challenging [10], in [1], we introduced a boundary model which describes very well the domain wall in the dual field theory, obtained through numerical relativity [3]. The model combines together the hydrodynamic description of the deconfined phase with a simple description of the confined phase, by adding an additional dynamical degree of freedom $\gamma$ describing the mixing of the two phases (see also [11, 12]). Specifically, the total energy-momentum tensor of the theory is a mixture of the hydrodynamic tensor of the deconfined phase and the trivial $T^{\mu\nu}=\eta^{\mu\nu}$ of the confined phase, together with a contribution describing the domain wall ($T^{\mu\nu}_{\Sigma}$):
@@
Boost invariant expansion. Boost invariant expansion, first introduced by Bjorken [17], can be seen as a way of modelling extremely energetic heavy-ion collisions, where the system undergoes expansion along the longitudinal (collision) axis. Consequently, this leads to decreasing energy density and hence also decreasing temperature. The inherent time-dependence and simplicity led boost-invariant expansion to be a fruitful arena for investigating various dynamical phenomena in holography (see e.g. [18, 19, 20, 21, 22, 23]). The expansion automatically drives the cooling system through any phase transitions or crossover to the low energy confined phase. Technically, one writes the standard Minkowski metric in adapted coordinates
@@
Neglecting the explicit temporal derivatives, the spatial part can be translated into a virial equation, but modulated by the scale factor $a(t)$. This indicates that the width of the dissolving blob shrinks with $a(t)$ in contrast to staying constant in the boost-invariant case. The virial theorem also leads to the conclusion of a rise of the temperature of the dissolving fluid. This is in very good agreement with the numerical simulations as shown in Fig. 9. Note that the observed behavior is driven by the expansion of the universe, and qualitatively different from heated droplets in earlier studies (see [28, 29] and references therein) in particular due to the rise of the temperature above $T_{c}$.